OpenStudy (anonymous):

solve for 0<_x<2pi 2sinxcosx=tanx

OpenStudy (freckles):

\[2 \sin(x) \cos(x)=\frac{sin(x)}{\cos(x)} \] \[2 \sin(x) \cos^2(x)=\sin(x)\] \[2 \sin(x) \cos^2(x)-\sin(x)=0\] \[\sin(x)[2 \cos^2(x)-1]=0\] set both factors equal to 0

OpenStudy (anonymous):

thanks

OpenStudy (freckles):

np